OpenStudy (anonymous):

Solve for x. 3x+12=9 Am I correct? would it be -1??

OpenStudy (anonymous):

That's right!

OpenStudy (anonymous):

alright thanks

OpenStudy (ranga):

x = -1 is correct. All you have to do is to put x = -1 in the equation and see if it is correct: 3(-1) + 12 = -3 + 12 = 9 which is the right hand side. So x = -1 is correct.

OpenStudy (anonymous):

oh yeah that makes sense. thank you